# On zeros of polynomials associated with Heun class equations

### v2

@@ Abstract.
Schäfke and Schmidt established that the asymptotics of the coefficients of the local solution to some linear differential equation is related to global structures of solutions. The Heun class equations have the accessory parameters, and we investigate the polynomials whose variable is the accessory parameter which appears as the coefficients of the local solution. By calculating the zeros of the polynomials numerically, we obtain the data of the spectral related to the Heun class equations numerically.
@@ 1. Introduction
Motivated by Lisovyy-Naidiuk [6] and Schäfke-Schmidt [11], we study Heun equations from the perturbative viewpoint by focusing to the accessory parameter. We now explain a result of Schäfke and Schmidt briefly in the situation of Heun’s differential equation. It is known by the theory of the regular singularity that a basis of the local solutions of the Heun equation about $z=0$ is obtained as
@@
The coefficients $d_{1}$ and $d_{2}$ are complex numbers. The situation such that $d_{2}=0$ is special, because the holomorphic solution about $z=0$ is continued to be a holomorphic solution about $z=1$, and it would be special in the connection problem related to application to physics. Schäfke and Schmidt [11] established the formula
@@
for the Heun equation under the condition $|t|>1$ and $\gamma,\delta\not\in\mathbb{Z}$. They established the formula as Theorem 2.1 in a more general setting. We can regard the values $d_{2}$ and $c_{k}$ as the functions on the accessory parameter $q$. By equation (1.7), it would be natural to expect that the position of the zeros of $d_{2}$ on the parameter $q$ resembles to the position of the zeros of $c_{k}$ for sufficiently large $k$.
In this paper, we investigate the zeros of the coefficient $c_{k}$ with respect to the accessory parameter $q$. It is shown that $c_{k}$ satisfies three-term recursive relations, and $c_{k}$ is a polynomial of degree $k$ on the variable $q$. We adopt a perturbative approach with respect to the variable $1/t$. In our approach, we obtain the hypergeometric equation as $1/t\to 0$. In the case $01/t=$ 0, the zeros of $c_{k}$ is calculated explicitly, and we try to understand the zeros of $c_{k}$ as a series expansion on the variable $1/t$. On the singly confluent Heun equation and the reduced singly confluent Heun equation, we can adopt similar approach, and we formulate by the unified notation.
On the other hand, we can calculate the value $c_{k}$ explicitly for small $k$ by computer, and also calculate the zeros of $c_{k}$ numerically. Then, we can observe the behaviours of the zeros as $k$ varies. We report them in Sections 5 and 6.
This paper is organized as follows. In Section 2, we review a remarkable result obtained in the paper [11] of Schäfke and Schmidt. In Section 3, we discuss three-term relations for coefficients $c_{k}$ of the local solution for Heun class equations with paying attention to the accessory parameter. In Section 4, we discuss perturbative expansions for zeros of $c_{k}$ with respect to the accessory parameter. In Section 5, we discuss the Lamé equation and numerical calculations on it. In Section 6, we discuss the Mathieu equation, the Whittaker-Hill equation and numerical calculations on them. In Section 7, we give concluding remarks.
@@ 2. Remarkable result of Schäfke and Schmidt
(2.6) $y_{01}=\eta_{0}(z)=\sum_{k=0}^{\infty}\frac{\tau^{0}_{k}}{\Gamma(1-\mu_{0}+k)}z^{k},\;y_{11}=\eta_{1}(z)=\sum_{k=0}^{\infty}\frac{\tau^{1}_{k}}{\Gamma(1-\mu_{1}+k)}(1-z)^{k},$
where $\tau^{0}_{0}=1=\tau^{1}_{0}$ and $\Gamma(\>\cdot\>)$ is the gamma function. Since the exponents about $z=1$ are 00 and $\mu_{1}$, we can take a basis of solutions to equation (2.1) as $y_{11}$ and $y_{12}$ where
$y_{12}=(1-z)^{\mu_{1}}\Big{(}\frac{1}{\Gamma(1+\mu_{1})}+\sum_{k=1}^{\infty}\frac{\tilde{\tau}^{1}_{k}}{\Gamma(1+\mu_{1}+k)}(1-z)^{k}\Big{)}.$
@@
be the Wronskian of two arbitrary solutions $y_{1}(z)$, $y_{2}(z)$ of equation (2.1). Then, there exists a unique constant $[y_{1}(z),y_{2}(z)]\in\mathbb{C}$ such that
[…]
@@
(2.9) $Q=[y_{01},y_{11}]=[\eta_{0}(z),\eta_{1}(z)].$
The elements of the connection matrix on Floquet solutions are expressed by using $Q$ ([11, Proposition 2.14]). We write a solution $y_{01}$ to equation (2.1) by using a basis $\{y_{11},y_{12}\}$ as
(2.10) $y_{01}=\tilde{d}_{1}y_{11}+\tilde{d}_{2}y_{12}.$
@@
(2.13) $y^{\langle 0\rangle}_{1}=\Gamma(1-\mu_{0})y_{01},\;y^{\langle 1\rangle}_{1}=\Gamma(1-\mu_{1})y_{11},\;y^{\langle 1\rangle}_{2}=\Gamma(1+\mu_{1})y_{12},$
@@ 3. Three-term relations for Heun class equations
Recall that Heun’s differential equation is given by
@@
By substituting it into equation (3.1) multiplied by $z(z-1)(z-t)$, we obtain the recursive relations
@@
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(i+1)(i+\gamma)tc_{i+1}=0,$
for $c_{i}$ $(i\geq 1)$. The solution $y$ is denoted by $Hl(t,q;\alpha,\beta,\gamma,\delta;z)$ in [10]. The coefficients $c_{i}$ are polynomial in the variable $q$, and the degree is equal to $i$. Set $q=Bt$, $t=1/s$ and
@@
(3.5) $\frac{d^{2}y}{dz^{2}}+\left(\frac{\gamma}{z}+\frac{\delta}{z-1}-\frac{s\epsilon}{1-sz}\right)\frac{dy}{dz}+\frac{B-s\alpha\beta z}{z(z-1)(1-sz)}y=0,$
@@
for each $B\in\mathbb{C}$, if $|s|<1$ and $\gamma,\delta\not\in\mathbb{Z}$.
If $s=0$, then equation (3.5) is in the form of the hypergeometric equation of Gauss, and the parameter $B$ is related with the local exponents about $z=\infty$. In this case, the values $d_{1}(B)$ and $d_{2}(B)$ can be calculated explicitly as the connection formula.
Recall that the singly confluent Heun equation is given by
(3.10) $\frac{d^{2}y}{dz^{2}}+\left(4p+\frac{\gamma}{z}+\frac{\delta}{z-1}\right)\frac{dy}{dz}+\frac{4p\alpha z-\sigma}{z(z-1)}y=0,$
We look into the solution written as
(3.11) $y=\sum_{i=0}^{\infty}c_{i}z^{i},\;c_{0}=1.$
By substituting it into equation (3.10) multiplied by $z(z-1)$, we obtain the recursive relations
(3.12) $4p(k-1+\alpha)c_{k-1}+[k(k-4p+\gamma+\delta-1)-\sigma]c_{k}-(k+1)(k+\gamma)c_{k+1}=0,$
where $c_{-1}=0$. See [10, pp. 101]. The solution $y$ is denoted by $Hc(p,\alpha,\gamma,\delta,\sigma;z)$ in [10]. The coefficients $c_{k}$ are polynomial in the variable $\sigma$, and the degree is equal to $k$.
Set $4p=-s$, $\sigma=-B$ and
@@
Then, the singly confluent Heun equation is written as
@@
Let $y^{\langle 1\rangle}_{1}$ and $y^{\langle 1\rangle}_{2}$ be a basis of the local solutions about $z=1$ which are written as equation (1.5). Then, the solution $y$ in equation (3.13) is written as
(3.17) $y=d_{1}(B)y^{\langle 1\rangle}_{1}+d_{2}(B)y^{\langle 1\rangle}_{2},$
and it follows from Theorem 2.1 that
for each $B,s\in\mathbb{C}$, if $\gamma,\delta\not\in\mathbb{Z}$.
Recall that the reduced singly confluent Heun equation was given in equation (1.3), and we change the parameters of it by $t=s$ and $\lambda=B$. Then, we obtain the differential equation
@@
Set
(3.20) $y=\sum_{k=0}^{\infty}c_{k}(B)z^{k}$
and substitute it into the reduced singly confluent Heun equation. Then, we have the recursive relation
(3.21) $c_{m+1}(B)=\frac{(B+D_{m})c_{m}(B)-sc_{m-1}(B)}{(m+1)(m+\gamma)},\;D_{m}=m(m-1+\gamma+\delta),$
where $c_{-1}(B)=0$. Note that this equation is also written as equation (3.15), where
@@ 4. Perturbative expansions for zeros
$m=0,1,2,\dots.$ If $s=0$, then we have
(4.2) $c_{m+1}(B)|_{s=0}=R_{m+1}(B+D_{0})(B+D_{1})\dots(B+D_{m}),\;R_{m+1}=\frac{1}{(m+1)!(\gamma)_{m+1}},$
where $(a)_{m+1}=a(a+1)\cdots(a+m)$.
@@
(4.3) $c_{m+1}(B)=c_{m+1}^{[0]}(B)+c_{m+1}^{[1]}(B)s+c_{m+1}^{[2]}(B)s^{2}+\dots,$
$c_{m+1}^{[0]}(B)=R_{m+1}(B+D_{0})(B+D_{1})\dots(B+D_{m}),$
where $c_{m+1}^{[j]}(B)$ are not dependent on $s$. The following proposition is shown immediately.
@@
Let $p$ be a non-negative integer. If $k=0,\cdots,p$, then we have
@@
We investigate the zeros of the polynomials $c_{m+1}(B)$ under the assumption that $D_{0},D_{1},D_{2},\dots$ are mutually different. We have $c_{m+1}(-D_{k})|_{s=0}=0$ for $k=0,1,\dots,m$. Since $D_{0},D_{1},D_{2},\dots$ are mutually different, we have $\frac{\partial}{\partial B}c_{m+1}(B)|_{s=0}=(D_{0}-D_{k})\dots(D_{k-1}-D_{k})(D_{k+1}-D_{k})\dots(D_{m}-D_{k})\neq 0$, and it follows from the analytic implicit function theorem that there exists a positive number $\varepsilon_{m+1}$ and the analytic functions $D_{k,m+1}(s)$ defined on the open disk $|s|<\varepsilon_{m+1}$ for $k=0,\dots,m+1$ such that $D_{k,m+1}(0)=D_{k}$ and
@@
Then, we have
(4.7) $0=c_{m+1}^{[0]}(-D_{k,m+1}(s))+c_{m+1}^{[1]}(-D_{k,m+1}(s))s+c_{m+1}^{[2]}(-D_{k,m+1}(s))s^{2}+\dots.$
We replace $D_{k}^{[j],(m+1)}$ to the indeterminant $x^{[j]}$ and we write
(4.8) $\sum_{j=0}^{\infty}c_{m+1}^{[j]}\Bigl{(}-D_{k}-\sum_{j^{\prime}=1}^{\infty}x^{[j^{\prime}]}s^{j^{\prime}}\Bigr{)}=\sum_{j=1}^{\infty}p_{j}s^{j}.$
Then, we have $p_{0}=0$ and $p_{j}$ is a polynomial of $x^{[1]},\dots,x^{[j]}$ such that the term containing $x^{[j]}$ is written as
(4.9) $-(D_{0}-D_{k})\dots(D_{k-1}-D_{k})(D_{k+1}-D_{k})\dots(D_{m}-D_{k})x^{[j]}.$
Therefore, if we impose the condition $p_{1}=0,p_{2}=0,\dots$, then the indeterminants $x^{[1]},x^{[2]},\dots$ are determined recursively, and it follows from equation (4.7) that $x^{[j]}=D_{k}^{[j],(m+1)}$ for $j=1,2,\dots$. We obtain the following proposition.
@@
Assume that $m\in\mathbb{Z}_{\geq 2}$. By substituting $B=-D_{m-1}+sA$ in the recursive equation (4.1) where $m$ is replaced with $m-1$ or $m-2$, we have
@@
Therefore, we obtain the proposition for the case $m\in\mathbb{Z}_{\geq 2}$. The case $m=1$ is shown similarly. ∎
@@
Therefore, the condition
(4.24) $A=-E_{m}-\frac{m(m+\gamma-1)}{D_{m}-D_{m-1}}F_{m}$
is equivalent to $c_{m+1}(B_{m}^{(m+1)})=O(s^{2})$. ∎
@@
(((([8])))) If $k\leq m-1$ and $m\leq m^{\prime}$, then we have $D_{k}^{[1],(m+1)}=D_{k}^{[1],(m^{\prime}+1)}$.
@@
By the definition of $D_{k}^{[1],(m+1)}$, we have $c_{m+1}(-D_{k}-D_{k}^{[1],(m+1)}s)=O(s^{2})$. It follows from Proposition 4.1 and the condition $k\leq m-1$ that $c_{m}(-D_{k}-As)=O(s)$ for any $A$, and we have $sF_{m+1}c_{m}(-D_{k}-D_{k}^{[1],(m+1)}s)=O(s^{2})$. Hence, we obtain $c_{m+2}(-D_{k}-D_{k}^{[1],(m+1)}s)=O(s^{2})$ by equation (4.25). On the other hand, it follows from the definition of $D_{k}^{[1],(m+2)}$ that $c_{m+2}(-D_{k}-D_{k}^{[1],(m+2)}s)=O(s^{2})$. Therefore, we have $D_{k}^{[1],(m+1)}=D_{k}^{[1],(m+2)}$ by applying Proposition 4.2. ∎
It follows from Proposition 4.5 that $D_{k}^{[1],(m+1)}=D_{k}^{[1],(k+2)}$, if $k\leq m-1$. By combining with Propositions 4.2, 4.3 and 4.4, we obtain the following theorem.
@@
(((([8])))) Assume that $D_{0},D_{1},D_{2},\dots$ are mutually different. Write
(4.26) $B_{k}^{(m+1)}=-D_{k}+\Bigl{\{}-E_{k}-\frac{k(k-1+\gamma)}{D_{k}-D_{k-1}}F_{k}-\frac{(k+1)(k+\gamma)}{D_{k}-D_{k+1}}F_{k+1}\Bigr{\}}s$
for $k=0,\cdots,m-1$, and
Then, $B=B_{k}^{(m+1)}$ represents the expansion of the solution to $c_{m+1}(B)=0$ in the variable $s$ up to the first degree for $k=0,\cdots,m$.
@@
Let $j$ be a positive integer. If $k+j\leq m$ and $m\leq m^{\prime}$, then we have $D_{k}^{[j],(m+1)}=D_{k}^{[j],(m^{\prime}+1)}$. If $m\geq k+j$, then we have $D_{k}^{[j],(m+1)}=D_{k}^{[j],(k+j+1)}$.
@@
It follows from Proposition 4.7 that the expansion of the zero of $c_{m+1}(B)=0$ labelled by $-D_{k}$ up to the term $s^{j}$ does not depend on $m$, if $j\leq m-k$. Namely, the expansion of the zero is written as
(4.29) $-D_{k,m+1}(s)=-D_{k}-\sum_{j^{\prime}=1}^{j}D_{k}^{[j^{\prime}],(k+j^{\prime}+1)}s^{j^{\prime}}+O(s^{j+1}),\quad c_{m+1}(-D_{k,m+1}(s))=0,$
for $j\leq m-k$. On the coefficient of $s^{2}$, we have the following proposition.
@@
+[$\quad+\Bigl{[}\Bigl{\{}\frac{(m-2)(m-3+\gamma)F_{m-2}}{(D_{m-2}-D_{m-3})^{2}}+\frac{(m-1)(m-2+\gamma)F_{m-1}}{(D_{m-2}-D_{m-1})^{2}}\Bigr{\}}$
@@
+[$\quad+\Bigl{[}\Bigl{\{}\frac{\gamma F_{1}}{(D_{1}-D_{0})^{2}}+\frac{2(1+\gamma)F_{2}}{(D_{1}-D_{2})^{2}}\Bigr{\}}\Bigl{\{}E_{1}+\frac{\gamma F_{1}}{D_{1}-D_{0}}+\frac{2(1+\gamma)F_{2}}{D_{1}-D_{2}}\Bigr{\}}$
@@
We obtain the value $A$ by a straightforward calculation, and we have equation (4.30). The cases $m=2,3$ are shown similarly. ∎

### v3

@@ Abstract.
The Heun class equations have the accessory parameters, and we investigate the polynomials, whose variable is the accessory parameter, which appear as the coefficients of a local solution. Our study is motivated by the result of Schäfke and Schmidt that the asymptotics of the coefficients of a local solution to some linear differential equation is related to global structures of solutions. The zeros of the polynomials are related to the spectral problem of the Heun class equation. We obtain exact results on the zeros from a perturbative approach, and we show some tables of the zeros by numerical calculations, which support our strategy.
@@ 1. Introduction
Motivated by Lisovyy-Naidiuk [6] and Schäfke-Schmidt [11], we study Heun equations from the perturbative viewpoint by focusing to the accessory parameter. We now explain a result of Schäfke and Schmidt briefly in the situation of Heun’s differential equation. It is known by the theory of the regular singularity that a basis of the local solutions of the Heun equation (1.1) about $z=0$ is obtained as
@@
The coefficients $d_{1}$ and $d_{2}$ are complex numbers. The situation such that $d_{2}=0$ is special, because the holomorphic solution about $z=0$ is continued to be a holomorphic solution about $z=1$, and it would be special in the connection problem related to application to physics. In this case, the solution to the Heun equation which is simultaneously holomorphic about $z=0$ and $1$ is called the Heun function. By the way, Schäfke and Schmidt [11] established the formula
@@
for the Heun equation under the condition $|t|>1$ and $\gamma,\delta\not\in\mathbb{Z}$. They established the formula as Theorem 2.1 in a more general setting. We can regard the values $d_{2}$ and $c_{k}$ as the functions on the accessory parameter $q$. By equation (1.7), it would be natural to expect that parts of the position of the zeros of $d_{2}$ on the parameter $q$ resemble to parts of the position of the zeros of $c_{k}$ for sufficiently large $k$.
In this paper, we replace the accessory parameter to the parameter $B$, and investigate the zeros of the coefficient $c_{k}(=c_{k}(B))$ with respect to the parameter $B$. It is shown that $c_{k}(B)$ is a polynomial of degree $k$ on the variable $B$. We set $s=1/t$ and adopt a perturbative approach with respect to the variable $s$. In our approach, we obtain the hypergeometric equation as $s\to 0$. In the case $s=0$, the zeros of $c_{k}(B)$ are calculated explicitly, and we try to understand the zeros of $c_{k}(B)$ as a series expansion on the variable $s$. On the singly confluent Heun equation and the reduced singly confluent Heun equation, we can adopt similar approachs, and we formulate them by the unified notation.
On the other hand, we can calculate the polynomials $c_{k}(B)$ explicitly for small $k$ by computer, and also calculate the zeros of $c_{k}(B)$ numerically. Then, we can observe the behaviours of the zeros as $k$ varies. We report them in Sections 5 and 6.
This paper is organized as follows. In Section 2, we review a remarkable result obtained in the paper [11] of Schäfke and Schmidt. In Section 3, we discuss three-term relations for coefficients of the local solution for Heun class equations with paying attention to the accessory parameter. In Section 4, we discuss perturbative expansions for zeros of $c_{k}(B)$ with respect to the accessory parameter. The principal results are Proposition 4.7 and Theorems 4.6 and 4.9. In Section 5, we discuss the Lamé equation and numerical calculations on it. In Section 6, we discuss the Mathieu equation, the Whittaker-Hill equation and numerical calculations on them. In Section 7, we give concluding remarks.
@@ 2. Remarkable result of Schäfke and Schmidt
(2.6) $y_{01}(z)=\sum_{k=0}^{\infty}\frac{\tau^{0}_{k}}{\Gamma(1-\mu_{0}+k)}z^{k},\quad y_{11}(z)=\sum_{k=0}^{\infty}\frac{\tau^{1}_{k}}{\Gamma(1-\mu_{1}+k)}(1-z)^{k},$
where $\tau^{0}_{0}=1=\tau^{1}_{0}$ and $\Gamma(\>\cdot\>)$ is the gamma function. Since the exponents about $z=1$ are 0 and $\mu_{1}$, we can take a basis of solutions to equation (2.1) as $y_{11}(z)$ and $y_{12}(z)$ where
$y_{12}(z)=(1-z)^{\mu_{1}}\Big(\frac{1}{\Gamma(1+\mu_{1})}+\sum_{k=1}^{\infty}\frac{\tilde{\tau}^{1}_{k}}{\Gamma(1+\mu_{1}+k)}(1-z)^{k}\Big).$
@@
be the Wronskian of two arbitrary solutions $y_{1}(z)$, $y_{2}(z)$ of equation (2.1). Then, the Wronskian satisfies $w^{\prime}+\{(1-\mu_{0})/z+(1-\mu_{1})/(z-1)+a(z)\}w=0$, and there exists a unique constant $y_{1}(z),y_{2}(z)]\in\mathbb{C}$ such that
(2.8) $w[y_{1},y_{2}](z)=[y_{1}(z),y_{2}(z)]z^{\mu_{0}-1}(1-z)^{\mu_{1}-1}\exp\Big(-\int_{1}^{z}a(\zeta)d\zeta\Big).$
@@
(2.9) $Q=[y_{01}(z),y_{11}(z)].$
The elements of the connection matrix on Floquet solutions are expressed by using $Q$ ([11, Proposition 2.14]). We write a solution $y_{01}(z)$ to equation (2.1) by using a basis $\{y_{11}(z),y_{12}(z)\}$ as
(2.10) $y_{01}(z)=\tilde{d}_{1}y_{11}(z)+\tilde{d}_{2}y_{12}(z).$
@@
(2.13) $y^{\langle 0\rangle}_{1}=\Gamma(1-\mu_{0})y_{01}(z),\;y^{\langle 1\rangle}_{1}=\Gamma(1-\mu_{1})y_{11}(z),\;y^{\langle 1\rangle}_{2}=\Gamma(1+\mu_{1})y_{12}(z),$
@@ 3. Three-term relations for Heun class equations
Recall that Heun’s differential equation was given by
@@
By substituting it into equation (3.1) multiplied by z(z1)(zt)z(z-1)(z-t), we obtain the recursive relations
@@
$\qquad\qquad\qquad\qquad\qquad\qquad+(i+1)(i+\gamma)tc_{i+1}=0,\;i=1,2,3,\dots.$
The solution $y$ in equation (3.2) was denoted by $Hl(t,q;\alpha,\beta,\gamma,\delta;z)$ in [10], and it was used to express other local solutions. The coefficients $c_{i}$ are polynomial in the variable $q$, and the degree is equal to $i$. Set $q=Bt$, $t=1/s$ and
@@
(3.5) $\frac{d^{2}y}{dz^{2}}+\left(\frac{\gamma}{z}+\frac{\delta}{z-1}-\frac{s\epsilon}{1-sz}\right)\frac{dy}{dz}+\frac{B-s\alpha\beta z}{z(z-1)(1-sz)}y=0.$
@@
for each $B\in\mathbb{C}$, if $|s|<1$ and $\gamma,\delta\not\in\mathbb{Z}$. We may expect that each position of the zeros of $d_{2}(B)$ is approximated by the position of the zeros of $c_{k}(B)$ as $k\to\infty$.
If $s=0$, then equation (3.5) is in the form of the hypergeometric equation of Gauss, and the parameter $B$ is related with the local exponents about $z=\infty$. In this case, the values $d_{1}(B)$ and $d_{2}(B)$ can be calculated explicitly by the connection formula.
Recall that the singly confluent Heun equation was given by
The solution written as equation (3.2) was denoted by $Hc(p,\alpha,\gamma,\delta,\sigma;z)$ in [10], and it was used to express other local solutions. We now change the parameters by 4p=s4p=-s and $\sigma=-B$. Then, the singly confluent Heun equation is written as
Set
@@
Let $y^{\langle 1\rangle}_{1}$ and $y^{\langle 1\rangle}_{2}$ be a basis of the local solutions about $z=1$ which are written as equation (1.5). Then, the solution $y$ in equation (3.12) is written as equation (3.8) and it follows from Theorem 2.1 that equation (3.9) holds for each $B,s\in\mathbb{C}$, if $\gamma,\delta\not\in\mathbb{Z}$.
The reduced singly confluent Heun equation was given in equation (1.3), and we change the parameters of it by $t=s$ and $\lambda=B$. Then, we obtain the differential equation
@@
Write a local solution $y$ as equation (3.12) and substitute it into the reduced singly confluent Heun equation. Then, we have the recursive relation
(3.16) $c_{m+1}(B)=\frac{\{B+m(m-1+\gamma+\delta)\}c_{m}(B)-sc_{m-1}(B)}{(m+1)(m+\gamma)},\;m=0,1,2,\dots,$
where $c_{-1}(B)=0$. Note that this equation is also written as equation (3.13) where
@@ 4. Perturbative expansions for zeros
(4.1) $c_{m+1}(B)=\frac{(B+D_{m}+sE_{m})c_{m}(B)-sF_{m}c_{m-1}(B)}{(m+1)(m+\gamma)},\;m=0,1,2,\dots.$
If $s=0$, then we have
(4.2) $c_{m+1}(B)|_{s=0}=R_{m+1}(B+D_{0})(B+D_{1})\cdots(B+D_{m}),\;R_{m+1}=\frac{1}{(m+1)!(\gamma)_{m+1}},$
where $(a)_{m+1}=a(a+1)\cdots(a+m)$. Therefore, the zeros of $c_{m+1}(B)|_{s=0}$ are $D_{0},-D_{1},$ $\dots,-D_{m}$, and we try to investigate the zeros of $c_{m+1}(B)$ as power series of the variable $s$.
@@
(4.3) $c_{m+1}(B)=R_{m+1}(B+D_{0})(B+D_{1})\cdots(B+D_{m})+d_{m+1}(B,s)s,$
where $d_{m+1}(B,s)$ is a polynomial in $B$ and $s$. The following proposition is shown immediately.
@@
Let $p$ be a non-negative integer. If $k\in\{0,1,\dots,p\}$, then we have
@@
We investigate the zeros of the polynomials $c_{m+1}(B)$ under the assumption that $D_{0},D_{1},D_{2},\dots$ are mutually different, which is equivalent to $\gamma+\delta\not\in\mathbb{Z}_{\leq 0}$ for $D_{m}=m(m-1+\gamma+\delta)$. In the propositions and the theorems below in this section, we also assume that $D_{0},D_{1},D_{2},\dots$ are mutually different. Let $k\in\{0,1,\dots,m\}$. We have $c_{m+1}(-D_{k})|_{s=0}=0$ and $\frac{\partial}{\partial B}c_{m+1}(B)|_{B=-D_{k},s=0}=R_{m+1}(D_{0}-D_{k})\cdots(D_{k-1}-D_{k})(D_{k+1}-D_{k})\cdots(D_{m}-D_{k})\neq 0$. Hence, it follows from the analytic implicit function theorem that there exists a positive number $\varepsilon$ and an analytic function $D_{k,m+1}(s)$ defined on the open disk $|s|<\varepsilon$ such that $D_{k,m+1}(0)=D_{k}$ and
@@
We discuss a characterization of the coefficients $D_{k}^{[j],(m+1)}$. We replace $D_{k}^{[j],(m+1)}$ to the indeterminant $x^{[j]}$ and we write
(4.7) $c_{m+1}\Bigl(-D_{k}-\sum_{j^{\prime}=1}^{\infty}x^{[j^{\prime}]}s^{j^{\prime}}\Bigr)=\sum_{j=0}^{\infty}p_{j}s^{j}.$
Then, we have $p_{0}=0$ and $p_{j}$ $(j\geq 1)$ is a polynomial of $x^{[1]},\dots,x^{[j]}$ such that the term containing $x^{[j]}$ is written as
(4.8) $-R_{m+1}(D_{0}-D_{k})\cdots(D_{k-1}-D_{k})(D_{k+1}-D_{k})\cdots(D_{m}-D_{k})x^{[j]},$
which follows from equation (4.3). Therefore, if we impose the condition $p_{1}=0,p_{2}=0,\dots$, then the indeterminants $x^{[1]},x^{[2]},\dots$ are determined recursively, and it follows from equation (4.5) that $x^{[j]}=D_{k}^{[j],(m+1)}$ for $j=1,2,\dots$. Thus, we obtain the following proposition.
@@
Assume that $m\in\mathbb{Z}_{\geq 2}$. By substituting $B=-D_{m-1}+sA$ in the recursive equation (4.1) where $m$ is replaced with m1m-1 or m2m-2, we have
@@
Therefore, we obtain the proposition for the case $m\in\mathbb{Z}_{\geq 2}$. The case $m=1$ is shown directly. ∎
@@
Therefore, if
(4.23) $A=-E_{m}-\frac{m(m+\gamma-1)}{D_{m}-D_{m-1}}F_{m},$
then we have $c_{m+1}(B_{m}^{(m+1)})=O(s^{2})$. ∎
@@
(([8])) If $k\leq m-1$ and $m\leq m^{\prime}$, then we have $D_{k}^{[1],(m+1)}=D_{k}^{[1],(m^{\prime}+1)}$.
@@
By Proposition 4.2, we have $c_{m+1}(-D_{k}-D_{k}^{[1],(m+1)}s)=O(s^{2})$. It follows from Proposition 4.1 and the condition $k\leq m-1$ that $c_{m}(-D_{k}-As)=O(s)$ for any AA, and we have $sF_{m+1}c_{m}(-D_{k}-D_{k}^{[1],(m+1)}s)=O(s^{2})$. Hence, we obtain $c_{m+2}(-D_{k}-D_{k}^{[1],(m+1)}s)=O(s^{2})$ by equation (4.24). Therefore, we have $D_{k}^{[1],(m+1)}=D_{k}^{[1],(m+2)}$ by applying Proposition 4.2. ∎
It follows from Proposition 4.5 that $D_{k}^{[1],(m+1)}=D_{k}^{[1],(k+2)}$, if $m\geq k+1$. By combining with Propositions 4.2, 4.3 and 4.4, we obtain the following theorem.
@@
(([8])) On the solution to $c_{m+1}(B)=0$ expressed as
(4.25) $B=-D_{k}-D_{k}^{[1],(m+1)}s+O(s^{2}),\quad k=0,1,\dots,m,$
we have
(4.26) $D_{k}^{[1],(m+1)}=D_{k}^{[1],(k+2)}=E_{k}+\frac{k(k-1+\gamma)}{D_{k}-D_{k-1}}F_{k}+\frac{(k+1)(k+\gamma)}{D_{k}-D_{k+1}}F_{k+1}$
for $k=1,\dots,m-1$ and
(4.27) $D_{0}^{[1],(m+1)}=E_{0}+\frac{\gamma}{D_{0}-D_{1}}F_{1},\quad D_{m}^{[1],(m+1)}=E_{m}+\frac{m(m-1+\gamma)}{D_{m}-D_{m-1}}F_{m}.$
@@
Let $j$ be a positive integer. If $k+j\leq m\leq m^{\prime}$, then we have $D_{k}^{[j],(m+1)}=D_{k}^{[j],(m^{\prime}+1)}$. In particular, we have $D_{k}^{[j],(m+1)}=D_{k}^{[j],(k+j+1)}$ under the condition $m\geq k+j$.
@@
It follows from Proposition 4.7 that the expansion of the zero of $c_{m+1}(B)$ labelled by $D_{k}$ up to the term $s^{j}$ does not depend on $m$, if $j\leq m-k$. Namely, the expansion is written as
(4.29) $-D_{k}-D_{k}^{[1],(k+2)}s-D_{k}^{[2],(k+3)}s^{2}-\cdots-D_{k}^{[j],(k+j+1)}s^{j}+O(s^{j+1})$
for $j\leq m-k$. Recall that the function $d_{2}(B)$ was defined by equation (3.8) and it is related to $c_{m+1}(B)$ by equation (3.9). We may expect that some of the zeros $d_{2}(B)$ are very close to some of the zeros of $c_{m+1}(B)$, if $m$ is large, and we propose the following conjecture.
Conjecture 1.
On the Heun equation, the confluent Heun equation and the reduced confluent Heun equation, the zeros of $d_{2}(B)$ are expanded as
(4.30) $-D_{k}-\sum_{j=1}^{\infty}D_{k}^{[j],(k+j+1)}s^{j}$
for $k=0,1,2,\dots$.
We now discuss the explicit expression of $D_{k}^{[2],(m+1)}$.
@@
$\quad+\Bigl[\Bigl\{\frac{(m-2)(m-3+\gamma)F_{m-2}}{(D_{m-2}-D_{m-3})^{2}}+\frac{(m-1)(m-2+\gamma)F_{m-1}}{(D_{m-2}-D_{m-1})^{2}}\Bigr\}$
@@
$\quad+\Bigl[\Bigl\{\frac{\gamma F_{1}}{(D_{1}-D_{0})^{2}}+\frac{2(1+\gamma)F_{2}}{(D_{1}-D_{2})^{2}}\Bigr\}\Bigl\{E_{1}+\frac{\gamma F_{1}}{D_{1}-D_{0}}+\frac{2(1+\gamma)F_{2}}{D_{1}-D_{2}}\Bigr\}$
@@
We obtain the value $A$ by a straightforward calculation, and we have equation (4.31). The cases $m=2,3$ are shown directly. ∎